# Discovering Effective Policies for Land-Use Planning with Neuroevolution

### v5

@@ 5.4. Interactive Evaluation
As an example, Figure 7 shows the grid cells from which the user has selected one. The system recommends a change in land use that results in an 18.09 tC/ha reduction in carbon emissions with a 15.77% land-use change. Using the sliders, the user can then explore alternative tradeoffs and modify specific elements of each suggestion. For each variation, the user can then obtain the estimated carbon reduction and land-use change percentage, and in this manner gain an understanding of the possible attainable solutions and confidence in the preferred choice.
6. Future Work
In the project so far we have not distinguished between the different types of crops and only partly considered regional variations in how land-use changes affect carbon stocks in soil and vegetation (by training different models for different regions). The assumed values for carbon stocks in soil and vegetation represent conditions of the recent past (Hansis et al., 2015); they do not consider environmental effects such as disturbances due to fires or droughts or better vegetation growth due to CO2 effects or a warmer climate, which can substantially impact terrestrial CO2 fluxes (e.g. Aragão et al., 2018; Pongratz et al., 2021; Rosan et al., 2024). Further, estimates of CO2 fluxes from land-use change also depend on the spatial resolution of the land-use change dataset (Ganzenmüller et al., 2022; Bultan et al., 2022). Extending the project with such details should lead to more precise and more fine-scale estimates of the impact of land-use change decisions in future work.
@@ 6. Future Work
The predictor creates point estimates of ELUC. A technique called Residual Input-Output kernel (RIO; Qiu et al., 2020) can be used to estimate confidence in these predictions. The method builds a Gaussian Process model of the residual error, quantifying the uncertainty. RIO is part of the Project Resilience codebase but was not yet used in the study in this paper.
Project Resilience also provides an alternative approach to prescription: Representing the prescriptor as a set of rules (Shahrzad et al., 2022). This approach has the advantage that the behavior is transparent and explainable, as demonstrated in prior experiments in other tasks (including NPI optimization (Miikkulainen et al., 2021)). Rule-set evolution on land-use optimization will be evaluated in the future.
To make the system more immediately useful to decision-makers, the cost metric could be refined to take into account region-specific costs of changing each particular land-use type to another. The approach can also be extended with preferences and further objectives. For instance, the decision maker could choose which crops should be preferred, which may be more actionable than the large-scale changes in the current dataset. It may be possible to optimize additional objectives such as minimizing nitrate pollution and water consumption, and maintaining or increasing food production. Similarly, instead of planning for a single year, it might be possible to develop prescriptors that recommend land-use changes for the next several years: For instance, what could be done in the next five years to reduce ELUC by a given target percentage. As data and simulations become more accurate in the future, such refined and extended decisions should become possible.
7. Conclusion
@@
Acknowledgements

### v7

@@ 5.6 Interactive Evaluation
As an example, Figure 9 shows the grid cells from which the user has selected one. The system recommends a change in land use that results in a 26.02 tC/ha reduction in carbon emissions with a 28.96% land-use change. Using the sliders, the user can then explore alternative tradeoffs and modify specific elements of each suggestion. For each variation, the user can then obtain the estimated carbon reduction and land-use change percentage, and in this manner gain an understanding of the possible attainable solutions and confidence in the preferred choice.
6 Future Work
The project so far has resulted in a promising proof-of-concept for utilizing ESP in land-use planning. In order to make it usable for real-world decision-makers, it needs to be extended with more fine-grained data. The models so far did not distinguish between the different types of crops and only partly considered regional variations in how land-use changes affect carbon stocks in soil and vegetation (through training different models for different regions). The assumed values for carbon stocks in soil and vegetation represent conditions of the recent past [14]; they do not consider environmental effects such as disturbances due to fires or droughts or better vegetation growth due to CO2 effects or a warmer climate, which can substantially impact terrestrial CO2 fluxes [e.g. 1, 27, 30]. Such additional features could be added as context input to the predictor and prescriptor models, and the actions could be updated accordingly to prescribe more detailed policies. Further, estimates of CO2 fluxes from land-use change also depend on the spatial resolution of the land-use change dataset [11, 2]. Extending the project with such details should lead to more precise and more fine-scale estimates of the impact of land-use change decisions, allowing decision-makers to utilize these dimensions in the real world.
@@ 6 Future Work
The predictor creates point estimates of ELUC. A technique called Residual Input-Output kernel [RIO; 28] can be used to estimate confidence in these predictions. The method builds a Gaussian Process model of the residual error, quantifying the uncertainty. RIO is part of the Project Resilience codebase but was not yet used in the study in this paper.
Project Resilience also provides an alternative approach to prescription: Representing the prescriptor as a set of rules [34]. This approach has the advantage that the behavior is transparent and explainable, as demonstrated in prior experiments in other tasks (including NPI optimization [22]). Rule-set evolution on land-use optimization will be evaluated in the future.
To make the system more immediately useful to decision-makers, the cost metric could be refined to take into account region-specific costs of changing each particular land-use type to another. The approach can also be extended with preferences and further objectives. For instance, the decision maker could choose which crops should be preferred, which may be more actionable than the large-scale changes in the current dataset. It may be possible to optimize additional objectives such as minimizing nitrate pollution and water consumption, and maintaining or increasing food production, similar to how minimizing crop change was optimized in Section 5.4. Further, instead of planning for a single year, it might be possible to develop prescriptors that recommend land-use changes for the next several years: For instance, what could be done in the next five years to reduce ELUC by a given target percentage. As data and simulations become more accurate in the future, such refined and extended decisions should become possible.
Even though land-use change is an important factor, it alone is not enough to reduce emissions to net zero. However, the framework designed in this project can be applied to other decision-making use cases as long as available context, action, and outcome data exist. A comprehensive such effort, covering energy policy and other factors, may eventually help reach climate-change goals.
7 Conclusion
@@
Appendix A Appendix: Using Location as a Proxy for Climate Information
Different locations have different emissions from land-use change due to the climate. They are affected by both the initial makeup of the land use and the conditions of the cell such as weather, soil, and type of vegetation. The BLUE model takes this type of vegetation into account as plant functional types (PFTs) when calculating ELUC. In order to test if the predictor model encodes such climate information, an experiment was run to see if similar trends arose in both the predicted ELUC from the NeuralNet and the carbon densities (i.e. potential for carbon release) computed from the BLUE model using plant functional types. Similarities in the results would indicate that the predictor model is able to learn climate information from plant functional types even when they are not explicitly passed in as features.
The carbon densities from the BLUE model are shown in Figure 10. To compute the corresponding values in the model, first note that crops have a significantly lower carbon density than vegetation (this effect can be seen e.g. in Figure 2, where conversion to crop from anything else results in a high ELUC). Thus, converting all land to crop would be similar to releasing all the carbon from the land, and the resulting ELUC can thus be seen as an estimate of its carbon density. To obtain this estimate, all land types were converted to crops globally across all data points in the year 2020, and the resulting ELUC predicted for this change using the NeuralNet. The result is plotted on a map in Figure 10.
Similar qualitative trends arise in both plots, with core primary forest regions like the Amazon and the Congo Basin having high carbon densities and predicted ELUCs. Secondary forest regions such as the eastern US, southern Canada, and southern China were picked up by the predictor as having a more moderate carbon density/ELUC. Additionally, drier regions like Central Asia tended towards lower carbon densities and ELUC values. The scales do not perfectly align, however it is more important to view overall trends rather than numeric similarities.
However, qualitatively higher ELUC values than carbon densities appeared in the Great Victoria Desert, northeastern Siberia, and northern Africa. This result was due to the fact that LUH2 classifies these regions as almost entirely primary nonforest vegetation, which the NeuralNet erroneously predicted to release a large amount of carbon. LUH2 uses country-level Food and Agriculture Organization (FAO) statistics when there is not enough data, and therefore defaults to a uniform distribution. While the predictor was able to pick up on trends in the other types of vegetation, it overestimated the ELUC for such regions.
To conclude, similar trends arise overall in carbon density estimates in BLUE and the model, indicating that the model captures climate information even when it is not explicitly given to it as features. Further alignment could be performed in the future using more specific data, however for the purposes of this proof-of-concept study, the amount of information encoded in the current features is sufficient.
Acknowledgments
@@
Funding Statement
This work received no specific grant from any funding agency, commercial or not-for-profit sectors.
Competing Interests
The authors report no competing interests.
Code and Data Availability Statement
Code and data for the experiments are available at https://github.com/Project-Resilience/mvp/tree/main/use_cases/eluc. Permanent snapshots of the data and code are archived on Zenodo at: 10.5281/zenodo.15226287 and 10.5281/zenodo.15231467
Author Contributions
Conceptualization: RM, OF, EM, JB, HC, BH; Data Curation: CS, OF, DY, JB, HC; Formal Analysis: EM, OF, DY; Investigation: RM, OF, DY, EM, BH; Methodology: RM, OF, DY, EM, BH; Project Administration: RM, OF, BH; Software: OF, EM, DY; Supervision: RM, OF, BH; Validation: RM, OF, EM, DY, BH; Visualization: RM, OF, EM, DY; Writing – Original Draft: RM, OF, DY, BH; Writing – Review & Editing: RM, OF, DY, EM, CS, BH.